OpenStudy (anonymous):

The correct answer to the integral (x² -1) from 1 to 3 is 6(2/3) I get the following: F(x) = x³ / 3 - x F(b) - F(a) = (3³/3-3) - (1³/3-1) = 6-(2/3)

hartnn (hartnn):

(1/3-1) = -2/3 6-(-2/3) = 6+2/3 = 6 2/3

OpenStudy (anonymous):

oh... of course... I thought 6(2/3) meant TIMES..... seemed a lot more strange than 6+2/3 Thanks again

hartnn (hartnn):

welcome ^_^